# Fine-Grained ECG-Text Contrastive Learning via Waveform Understanding Enhancement

### v1

@@ 3.2 Training Process
As discussed above, waveform features, which serve as intermediate products of diagnostic reasoning by doctors, are often not recorded in reports but are crucial for the model to learn ECG representations and understand the diagnostic process. Therefore, we propose a novel training process illustrated in Figure 2 consisting of three steps: (1) training the CLEP model using contrastive learning on original ECG-report pairs, (2) generating potential waveform features based on the original report using LLMs and validating them with CLEP, and (3) continuing to train the CLEP model with this augmented report containing waveform features to obtain the final FG-CLEP model.
@@
The key of our training process is to recover the missing waveform features in the report. Given a report, we query LLMs with the question, ‘What waveform features are most likely to be present in electrocardiograms with these symptoms?’ to identify potential overlooked waveform features. To format the results, we further instruct, ‘Organize these waveform features into a Python list, with each item representing a distinct waveform feature.’ Using this explicit chain-of-thought instruction Wei et al. (2022), we obtain a list of potential waveform features. However, LLM outputs are unreliable for two reasons: First, ECG waveform features and diagnoses are not in a one-to-one correspondence Jin (2018)—a single disease may present different waveform characteristics across individuals. Doctors can infer a diagnosis from waveform features, but not the other way around. Second, even if the non-bijective relationship is excluded, the LLM’s output is inherently unstable due to hallucination issues Huang et al. (2023); Günay et al. (2024).
@@
Thus, we validate LLM-generated waveform features by computing their similarity to the ECG signal using coarse CLEP, selecting only high-confidence waveform features for augmentation. These validated features are then incorporated into the original report for continued training of the Fine-Grained CLEP model. See Appendix B for pseudo code.
Table 1: Results of zero-shot classification. ENS: ensemble inference.
METS Li et al. (2024) 76.31 80.12 65.95 86.29 82.49 77.20
FG-METS 78.12$\uparrow$1.81 82.01$\uparrow$1.89 66.33$\uparrow$0.38 90.12$\uparrow$3.83 86.92$\uparrow$4.43 81.20$\uparrow$4.00
MERL Liu et al. (2024b) 74.20 75.70 65.90 78.50 82.80 74.40
FG-MERL 76.70$\uparrow$2.50 78.20$\uparrow$2.50 66.80$\uparrow$0.90 81.00$\uparrow$2.50 85.30$\uparrow$2.50 76.90$\uparrow$2.50
CLEP 77.50 81.85 66.29 88.60 85.15 80.10
FG-CLEP 79.28$\uparrow$1.78 83.57$\uparrow$1.72 67.77$\uparrow$1.48 92.31$\uparrow$3.71 88.24$\uparrow$3.09 82.46$\uparrow$2.36
CLEPENS 75.64$\downarrow$1.86 82.55$\uparrow$0.70 64.74$\downarrow$1.55 88.67$\uparrow$0.07 83.91$\downarrow$1.24 81.24$\uparrow$1.14
FG-CLEP ENS 79.68$\uparrow$0.40 83.65$\uparrow$0.08 70.79$\uparrow$3.02 91.52$\downarrow$0.79 87.08$\downarrow$1.16 84.60$\uparrow$2.14
@@ 4.1 Datasets
We pre-train the FG-CLEP framework using the MIMIC-ECG Gow et al. dataset and test it on the PTB-XL Wagner et al. (2020), CPSC2018 Liu et al. (2018), and CSN Zheng et al. (2022) datasets, following the benchmark proposed by Liu et al. (2024b). All the ECGs in the datasets are 12-lead recordings. The MIMIC-ECG dataset contains nearly 800,000 ECG-report pairs. To improve data quality, we excluded samples with an empty report or reports containing fewer than three words, removed reports without useful information, and discarded ECGs with unexpected situations. Details regarding the train:validation:test split and other dataset-specific information are provided in Appendix A.
@@
SimCLRChen et al. (2020) 63.41 69.77 73.53 60.84 68.27 73.39 54.98 56.97 62.52 51.41 69.44 77.73 59.78 68.52 76.54 59.02 67.26 73.20
BYOLGrill et al. (2020) 71.70 73.83 76.45 57.16 67.44 71.64 48.73 61.63 70.82 41.99 74.40 77.17 60.88 74.42 78.75 54.20 71.92 74.69
BarlowTwinsZbontar et al. (2021) 72.87 75.96 78.41 62.57 70.84 74.34 52.12 60.39 66.14 50.12 73.54 77.62 55.12 72.75 78.39 60.72 71.64 77.43
MoCo-v3Chen et al. (2021) 73.19 76.65 78.26 55.88 69.21 76.69 50.32 63.71 71.31 51.38 71.66 74.33 62.13 76.74 75.29 54.61 74.26 77.68
SimSiamChen and He (2021) 73.15 72.70 75.63 62.52 69.31 76.38 55.16 62.91 71.31 49.30 69.47 75.92 58.35 72.89 75.31 58.25 68.61 77.41
TS-TCCEldele et al. (2021) 70.73 75.88 78.91 53.54 66.98 77.87 48.04 61.79 71.18 43.34 69.48 78.23 57.07 73.62 78.72 55.26 68.48 76.79
CLOCSKiyasseh et al. (2021) 68.94 73.36 76.31 57.94 72.55 76.24 51.97 57.96 72.65 47.19 71.88 76.31 59.59 77.78 77.49 54.38 71.93 76.13
ASTCLWang et al. (2023) 72.51 77.31 81.02 61.86 68.77 76.51 44.14 60.93 66.99 52.38 71.98 76.05 57.90 77.01 79.51 56.40 70.87 75.79
CRTZhang et al. (2023) 69.68 78.24 77.24 61.98 70.82 78.67 46.41 59.49 68.73 47.44 73.52 74.41 58.01 76.43 82.03 56.21 73.70 78.80
ST-MEMNa et al. (2024) 61.12 66.87 71.36 54.12 57.86 63.59 55.71 59.99 66.07 51.12 65.44 74.85 56.69 63.32 70.39 59.77 66.87 71.36
MERLLiu et al. (2024b) 82.39 86.27 88.67 64.90 80.56 84.72 58.26 72.43 79.65 53.33 82.88 88.34 70.33 85.32 90.57 66.60 82.74 87.95
CLEP 84.04 88.79 89.82 69.09 86.08 92.50 67.89 72.35 82.59 61.79 91.86 90.18 83.12 93.42 96.56 63.00 80.03 93.35
FG-CLEP 84.89 89.51 90.77 69.96 85.75 92.62 68.91 74.80 85.42 68.99 91.35 94.08 83.35 93.60 96.65 62.59 79.35 93.46
4.2 Implementation Details
Pre-training Implementation: In the pre-training stage, we utilize a randomly initialized 1D-ResNet50 model He et al. (2016) as the ECG encoder and BioClinicalBERT Alsentzer et al. (2019) for text encoding. The AdamW optimizer is selected with a learning rate of $2\times 10^{-5}$ and a weight decay of $1\times 10^{-4}$. CLEP is pre-trained for 10 epochs with original reports and FG-CLEP is trained for another 3 epochs with fine-grained reports, using a cosine annealing scheduler for learning rate adjustments and a warmup phase for the first 10% of training steps. A batch size of 100 is maintained. The temperature parameters $t$ and $b$ are initialized to $\log 10$ and $-10$, respectively. The default hyperparameter $\lambda$ is set to 0.5 and the default threshold for selecting high-confidence waveform features is set to 0.95. We use LLaMA3-8B AI@Meta (2024) as our LLM to query potential waveform features and use vLLM Kwon et al. (2023) to speed up inference. All experiments used two NVIDIA A800 80GB GPUs, except LLaMA3-70B ablation, which used four.
@@ 4.3 Zero-Shot Ability
The zero-shot results are illustrated in Table 1. Both CLEP and FG-CLEP performed well. A detailed examination of the data reveals that FG-CLEP significantly outperforms CLEP on PTBXL-Form, PTBXL-Rhythm demonstrating that continue training using fine-grained reports substantially enhanced the model’s ability to capture local ECG waveform features. This improvement is particularly evident when using the ensemble method, which extends the label text to 12 leads (‘label in lead x’, where x represents any of the 12 leads). This further indicates FG-CLEP’s fine-grained waveform feature capture capability. However, the ensemble inference method often proves detrimental to CLEP, as seen in PTBXL-Super, PTBXL-Form, and CPSC2018. Additionally, we applied our generated fine-grained reports to other methods, METS Li et al. (2024) and MERL Liu et al. (2024b), to validate the generalizability. The results demonstrate that our fine-grained reports can also enhance the performance of these methods.
@@ 4.4 Linear Evaluation
We aim to evaluate the learned model transferability to downstream supervised tasks. We froze the ECG encoder and fine-tuned a randomly initialized linear classification head on the training data with binary cross-entropy loss. We compared a series of contrastive and generative self-supervised learning methods. Results in Table 2 show that FG-CLEP still achieves the best performances across all methods in most scenarios.
@@
5 Analysis
In this section, we visualize the semantic similarity matrix to identify false negatives, conduct an ablation study to validate FG-CLEP’s effectiveness, and assess its performance across various LLMs, text encoders, and ECG encoders. The reported metrics reflect the average zero-shot AUC across the aforementioned six datasets.
5.1 Semantic Similarity Matrix
We visualize the semantic similarity matrix in Figure 4. The left side shows the semantic similarity matrix from a random batch. As illustrated, ECGs and reports from different records may share similarities to some extent. Ignoring these similarities would result in a diagonal matrix with ones on the diagonal and zeros elsewhere, which is obviously wrong. The right side displays a semantic similarity matrix where the first 16 entries are normal ECGs and the last 16 are abnormal ECGs. The matrix effectively captures the semantic similarities of the normal ECGs.
Figure 4: The Heatmap of Semantic Similarity Matrix. Left: from a random batch; Right: with the first 16 as normal ECG and the last 16 as abnormal ECG.
5.2 Ablation Study
To evaluate the effectiveness of our proposed fine-grained pretraining strategy, we conducted an ablation study, with the results presented in Table 3. Specifically, we compared FG-CLEP with CLEP trained without the additional 3 epochs of fine-grained report training. The results demonstrate that FG-CLEP significantly outperforms CLEP. To further confirm that the performance gains stem from the fine-grained reports rather than additional training epochs, we trained CLEP for 3 extra epochs using the original reports. The findings indicate that CLEP converges within 10 epochs, and the additional training even risks overfitting, leading to a slight performance drop. This further validates the importance of fine-grained reports.
We also tested training FG-CLEP from scratch using fine-grained reports instead of continuing from CLEP. The results show no performance improvement but increased computational costs, supporting our default approach of continuing training from CLEP. Additionally, to verify the effectiveness of our model architecture in mitigating false negatives, we evaluated the performance using the original InfoNCE loss. The results reveal a significant performance drop, highlighting the efficacy of our proposed loss function in addressing false negatives.
@@ 5.2 Ablation Study
FG-CLEP (continued from CLEP) 82.27
FG-CLEP (trained from scratch) 82.28
CLEP (without fine-grained training) 79.92
CLEP (+3 epochs, original reports) 79.01
CLEP (InfoNCE loss only) 80.78
@@
Figure 5: Effect of Threshold and $\lambda$
5.3 Different Components
We conducted experiments to evaluate the performance of our framework using different LLMs, text encoders, and ECG encoders. The results are presented in Table 4. The findings indicate that our framework is robust across various components. Specifically, for different LLMs Abdin et al. (2024); Jiang et al. (2023a); AI@Meta (2024); Labrak et al. (2024); Ankit Pal (2024), larger LLM provide some performance improvements, though the gains are not substantial. While domain-specific LLMs possess more medical knowledge, our method requires formatting the waveform feature outputs, an area where domain-specific models are less effective, resulting in performance that does not surpass general-purpose models. For different text encoders Alsentzer et al. (2019); Gu et al. (2021); Jin et al. (2023); Lee et al. (2020), our framework consistently achieves significant improvements. Regarding ECG encoders He et al. (2016); Dosovitskiy et al. (2020), scaling up the model size leads to noticeable performance gains, and ResNet He et al. (2016) outperforms ViT Dosovitskiy et al. (2020) in capturing ECG features.

### v2

@@ 3.2 Coarse-to-Fine Training Process
Fine-grained alignment relies on fine-grained reports that include detailed waveform features. Although we achieve fine-grained alignment between tags and specific ECG segments, as discussed above, we observe that due to clinical habits, nearly half of the reports do not record waveform features as intermediate observations. Therefore, we propose a novel training process illustrated in Figure 2 consisting of three steps: (1) training the CLEP model using contrastive learning on original ECG-report pairs, (2) generating potential waveform features based on the original report using LLMs and validating them with CLEP, and (3) continuing to train the CLEP model with this augmented report containing waveform features to obtain the final FG-CLEP model.
@@
The key to our training process is to recover the missing waveform features in the report. Given a report, we query LLMs with the question, ‘What waveform features are most likely to be present in electrocardiograms with these symptoms?’ to identify potential overlooked waveform features. To format the results, we further instruct, ‘Organize these waveform features into a Python list, with each item representing a distinct waveform feature.’ Using this explicit chain-of-thought instruction (Wei et al., 2022), we obtain a list of potential waveform features. However, LLM outputs are unreliable for two reasons: First, ECG waveform features and diagnoses are not in a one-to-one correspondence (Jin, 2018)—a single disease may present different waveform characteristics across individuals. Doctors can infer a diagnosis from waveform features, but not the other way around. Second, even if the non-bijective relationship is excluded, the LLM’s output is inherently unstable due to hallucination issues (Huang et al., 2023; Günay et al., 2024).
@@
Thus, we validate LLM-generated waveform features by computing their similarity to the ECG signal using coarse CLEP, selecting only high-confidence waveform features for augmentation. These validated features are then incorporated into the original report for continued training of the Fine-Grained CLEP model.
3.3 False Negative Mitigation
To address the false negative problem exacerbated by tag-level fine-grained alignment, we introduce a semantic similarity matrix that computes the similarity between different tags. This matrix is incorporated into contrastive learning to guide the model in identifying and correcting false negatives.
Formally, the semantic similarity matrix $\mathbf{S}\in\mathbb{R}^{B\times B}$ is defined as follows.
$\mathbf{S}_{ij}=\text{sim}(\mathbf{e}_{i},\mathbf{e}_{j})\in[0,1].$ (8)
We integrate the semantic similarity matrix into the loss function to guide contrastive learning, following the approach in (Sun et al., 2023). This matrix captures the semantic similarity between tags from different reports, enabling the model to identify and correct false negative samples by ensuring that similar tags are aligned more effectively, even if they belong to different reports.
The loss term $L_{\text{fnm}}$ is defined as:
$L_{\text{fnm}}=\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{B}\left|\text{sim}(\mathbf{e}_{i},\mathbf{z}_{j})-\mathbf{S}_{ij}\right|.$ (9)
where $B$ is the batch size and $|\cdot|$ denotes the L1 distance.
The final loss function for FG-CLEP is the combination of the contrastive loss and the false negative mitigation loss:
$L=L_{\text{con}}+\lambda L_{\text{fnm}}.$ (10)
@@ 4.1 Datasets
We pre-train the FG-CLEP framework using the MIMIC-ECG (Gow et al., ) dataset and test it on the PTB-XL (Wagner et al., 2020), CPSC2018 (Liu et al., 2018), and CSN (Zheng et al., 2022) datasets, following the benchmark proposed by (Liu et al., 2024b). All the ECGs in the datasets are 12-lead recordings. The MIMIC-ECG dataset contains nearly 800,000 ECG-report pairs. To improve data quality, we excluded samples with an empty report or reports containing fewer than three words, removed reports without useful information, and discarded ECGs with unexpected situations. Details regarding the train:validation:test split and other dataset-specific information are provided in the Appendix.
4.2 Implementation Details
Pre-training Implementation: In the pre-training stage, we utilize a randomly initialized ViT model (Dosovitskiy et al., 2020) as the ECG encoder and BioClinicalBERT (Alsentzer et al., 2019) for text encoding. The whole ECG is divided into 60 non-overlapping patches. The AdamW optimizer is selected with a learning rate of $2\times 10^{-5}$ and a weight decay of $1\times 10^{-4}$. CLEP is pre-trained for 10 epochs with original reports and FG-CLEP is trained for another 3 epochs with fine-grained reports, using a cosine annealing scheduler for learning rate adjustments and a warmup phase for the first 10% of training steps. A batch size of 100 is maintained. The temperature parameters $t$ are initialized to $\log 10$. The default hyperparameter $\lambda$ is set to 0.5 and the default threshold for selecting high-confidence waveform features is set to 0.95. We use LLaMA3-8B (AI@Meta, 2024) as our LLM to query potential waveform features and use vLLM (Kwon et al., 2023) to speed up inference. All experiments used two NVIDIA A800 80GB GPUs, except LLaMA3-70B ablation, which used four.
Table 1: Results of zero-shot classification. ENS: ensemble inference. FG-: trained with fine-grained reports.
METS (Li et al., 2024) 76.31 80.12 65.95 86.29 82.49 77.20
[…]
MERL (Liu et al., 2024b) 74.20 75.70 65.90 78.50 82.80 74.40
[…]
CLEP 78.01 82.41 67.96 89.48 85.94 80.88
[…]
@@ 4.2 Implementation Details
SimCLR 63.41 69.77 73.53 60.84 68.27 73.39 54.98 56.97 62.52 51.41 69.44 77.73 59.78 68.52 76.54 59.02 67.26 73.20
BYOL 71.70 73.83 76.45 57.16 67.44 71.64 48.73 61.63 70.82 41.99 74.40 77.17 60.88 74.42 78.75 54.20 71.92 74.69
BarlowTwins 72.87 75.96 78.41 62.57 70.84 74.34 52.12 60.39 66.14 50.12 73.54 77.62 55.12 72.75 78.39 60.72 71.64 77.43
MoCo-v3 73.19 76.65 78.26 55.88 69.21 76.69 50.32 63.71 71.31 51.38 71.66 74.33 62.13 76.74 75.29 54.61 74.26 77.68
SimSiam 73.15 72.70 75.63 62.52 69.31 76.38 55.16 62.91 71.31 49.30 69.47 75.92 58.35 72.89 75.31 58.25 68.61 77.41
TS-TCC 70.73 75.88 78.91 53.54 66.98 77.87 48.04 61.79 71.18 43.34 69.48 78.23 57.07 73.62 78.72 55.26 68.48 76.79
CLOCS 68.94 73.36 76.31 57.94 72.55 76.24 51.97 57.96 72.65 47.19 71.88 76.31 59.59 77.78 77.49 54.38 71.93 76.13
ASTCL 72.51 77.31 81.02 61.86 68.77 76.51 44.14 60.93 66.99 52.38 71.98 76.05 57.90 77.01 79.51 56.40 70.87 75.79
CRT 69.68 78.24 77.24 61.98 70.82 78.67 46.41 59.49 68.73 47.44 73.52 74.41 58.01 76.43 82.03 56.21 73.70 78.80
ST-MEM 61.12 66.87 71.36 54.12 57.86 63.59 55.71 59.99 66.07 51.12 65.44 74.85 56.69 63.32 70.39 59.77 66.87 71.36
MERL 82.39 86.27 88.67 64.90 80.56 84.72 58.26 72.43 79.65 53.33 82.88 88.34 70.33 85.32 90.57 66.60 82.74 87.95
CLEP 84.73 89.45 90.24 69.61 86.39 92.86 68.64 73.23 83.27 62.39 92.80 90.81 83.79 94.02 97.22 63.54 80.76 94.01
FG-CLEP 85.49 90.34 91.33 70.86 86.46 93.36 69.53 75.53 86.26 69.61 92.11 94.64 84.08 94.33 97.42 63.32 80.01 94.17
@@ 4.3 Zero-Shot Ability
The zero-shot results are illustrated in Table 1. Both CLEP and FG-CLEP performed well. A detailed examination of the data reveals that FG-CLEP significantly outperforms CLEP on PTBXL-Form, PTBXL-Rhythm demonstrating that continue training using fine-grained reports substantially enhanced the model’s ability to capture local ECG waveform features. This improvement is particularly evident when using the ensemble method, which extends the label text to 12 leads (‘label in lead x’, where x represents any of the 12 leads). This further indicates FG-CLEP’s fine-grained waveform feature capture capability. However, the ensemble inference method often proves detrimental to CLEP, as seen in PTBXL-Super, PTBXL-Form, and CPSC2018.
Additionally, we applied our generated fine-grained reports to other methods, METS (Li et al., 2024) and MERL (Liu et al., 2024b), to validate the generalizability. The results demonstrate that our fine-grained reports can also enhance the performance of these methods.
@@ 4.4 Linear Evaluation
We aim to evaluate the learned model transferability to downstream supervised tasks. We froze the ECG encoder and fine-tuned a randomly initialized linear classification head on the training data with binary cross-entropy loss. We compared a series of contrastive and generative self-supervised learning methods. Results in Table 2 show that FG-CLEP still achieves the best performances across all methods (Chen et al., 2020; Grill et al., 2020; Zbontar et al., 2021; Chen et al., 2021; Chen & He, 2021; Eldele et al., 2021; Kiyasseh et al., 2021; Wang et al., 2023; Zhang et al., 2023; Na et al., 2024; Liu et al., 2024b) in most scenarios.
@@
FG-CLEP(default) 83.03
w/o Fine-Grained Alignment 82.27
w/o Fine-Grained Report 80.78
CLEP + 3 epochs (original reports) 80.03
FG-CLEP (trained from scratch) 83.01
w/o False Negative Mitigation 81.67
Figure 3: Effect of Threshold and $\lambda$.
Figure 4: Visual activation maps generated by FG-CLEP and by the model without fine-grained alignment, using the text prompt ‘RSR’ Pattern in V1-V3’.
5 Analysis
In this section, we conduct a series of experiments to provide an in-depth analysis of FG-CLEP. The reported metrics reflect the average zero-shot AUC across the six datasets described above.
5.1 Ablation Study
To evaluate the effectiveness of our proposed fine-grained alignment with tag-specific ECG segments, the coarse-to-fine training with fine-grained reports, and the false negative mitigation loss function, we conducted a series of ablation studies. The results are presented in Table 3.
Specifically, when removing the fine-grained alignment and performing contrastive learning only at the whole-report and ECG level, the performance drops significantly. We also compared FG-CLEP with CLEP trained without the additional 3 epochs of fine-grained report training. The results demonstrate that FG-CLEP significantly outperforms CLEP. To further confirm that the performance gains stem from the fine-grained reports rather than additional training epochs, we trained CLEP for 3 extra epochs using the original reports. The findings indicate that CLEP converges within 10 epochs, and the additional training even risks overfitting, leading to a slight performance drop. This further validates the importance of fine-grained reports. We also tested training FG-CLEP from scratch using fine-grained reports instead of continuing from CLEP. The results show no performance improvement but increased computational costs, supporting our default approach of continuing training from CLEP. Additionally, to verify the effectiveness of mitigating false negatives, we evaluated the performance without $L_{\text{fnm}}$. The results reveal a significant performance drop, highlighting the efficacy of our proposed loss function in addressing false negatives.
@@ 5.1 Ablation Study
5.2 Visual Activation Map
To further demonstrate the fine-grained alignment between report tags and specific ECG segments, we present visual activation maps generated by FG-CLEP and by the model without fine-grained alignment using the text prompt ‘RSR’ Pattern in V1-V3’, which is a key waveform feature for diagnosing right bundle branch block. As shown in Figure 5, FG-CLEP effectively captures ECG segments relevant to the text prompt, whereas the model without fine-grained alignment fails to accurately localize the specific ECG segments. This further validates the effectiveness of our approach in achieving fine-grained alignment.
5.3 Validation of Fine-Grained Report
In our training dataset MIMIC-ECG, nearly 50% of reports do not mention any waveform features. As mentioned earlier, rather than directly using the waveform features generated by the LLM as ground truth, we validate these features with CLEP to produce fine-grained reports. To evaluate the accuracy and reliability of the generated fine-grained reports, as well as the effectiveness of CLEP for validation, we randomly selected 100 ECGs that lacked waveform features. Three medical students independently annotated these ECGs for five key waveform features. The final labels were determined by majority vote, and the accuracy of the generated fine-grained reports was calculated accordingly. Since the same waveform feature may have different (granular) expressions, we computed the similarity between the generated reports and the five waveform features to obtain AUC values. As shown in Table 4, the AUC of waveform features directly generated by the LLM is relatively low, which aligns with expectations due to the hallucination issues of LLMs and the non-bijective relationship between ECG waveform features and diagnoses. However, after validation with CLEP, the AUC of the fine-grained reports improves significantly, demonstrating the effectiveness and reliability of the coarse-to-fine training process in generating fine-grained reports.
Table 4: AUC of Generated Fine-Grained Reports (with/without verification).
Waveform Feature w/o Veri. w Veri.
Non-specific ST elevation 71.42 83.42
Long QT-interval 84.25 93.25
Abnormal QRS 76.68 91.68
Prolonged PR interval 75.13 86.13
Inverted T-waves 74.79 88.79
Figure 5: The Heatmap of Semantic Similarity Matrix. Left: a random batch; Right: with the first 16 as normal ECG and the last 16 as abnormal ECG.
5.4 Semantic Similarity Matrix
We visualize the semantic similarity matrix in Figure 5. The left side shows the semantic similarity matrix from a random batch. As illustrated, ECGs and tags from different records may share similarities to some extent. Ignoring these similarities would result in a diagonal matrix with ones on the diagonal and zeros elsewhere, which is obviously wrong. The right side displays a semantic similarity matrix where the first 16 entries are normal ECGs and the last 16 are abnormal ECGs. The matrix effectively captures the semantic similarities of the normal ECGs.